(c) Suppose n = 50 and p = 0.19. (For each answer, enter a number. Use 2 decimal places.)
n·p =
9.50

Correct: Your answer is correct.

n·q =
40.50

Correct: Your answer is correct.


Can we approximate p̂ by a normal distribution? Why? (Fill in the blank. There are four answer blanks. A blank is represented by _____.)

_____, p̂ _____ be approximated by a normal random variable because _____ _____.

first blank
Yes
No
Correct: Your answer is correct.
second blank
can
cannot
Correct: Your answer is correct.
third blank
n·q exceeds
n·p exceeds
n·q does not exceed
both n·p and n·q exceed
n·p and n·q do not exceed
n·p does not exceed
Correct: Your answer is correct.
fourth blank (Enter an exact number.)


What are the values of μp̂ and σp̂? (For each answer, enter a number. Use 3 decimal places.)
μp̂ = mu sub p hat =


σp̂ = sigma sub p hat =
1
Expert's answer
2020-01-06T06:10:13-0500

(c) Suppose n = 50 and p = 0.19.


"n\cdot p=50\cdot 0.19=9.50""n\cdot p=9.50"

"n\cdot q=50\cdot (1-0.19)=40.50""n\cdot q=40.50"

Can we approximate p̂ by a normal distribution? Why?

Yes, p̂ can be approximated by a normal random variable because both n·p and n·q exceed 5.


What are the values of μp̂ and σp̂?

"\mu_ {\hat{p}}=np=9.500"

"\sigma_ {\hat{p}}=\sqrt{npq}=2.774"